# Inductive Gradient Adjustment for Spectral bias in implicit neural representations

### v1

@@ A.2 Basic Lemmas
where $\Delta f({\bm{x}}_{i},\Theta_{t})=({\bm{r}}_{t+1}-{\bm{r}}_{t})_{i}$ denotes the dynamics of $f({\bm{x}},\Theta)$ at time step $t$; $R_{0},\kappa$ are constants; $\nabla_{\Theta_{t}}f({\bm{x}}^{j_{i}},\Theta)^{\top}\nabla_{\Theta_{t}}f({\bm{x}}^{j},\Theta)$ is the entry in ${\tilde{{\bm{K}}}_{e}}(i,j)$.
@@
From the equation 18 of lemma 4, we consider the following error:
@@
$\leq\epsilon\|\nabla_{\Theta_{t}}f({\bm{X}},\Theta_{t}){\bm{I}}{\bm{r}}_{t}\|\leq\epsilon\frac{\kappa R_{0}}{\sqrt{m}}$
@@
We define that $\epsilon^{\prime}=A+B$. By the lemma 4, we have that:
@@
${\bm{r}}^{i}_{t+1}=p\tilde{{\bm{K}}}_{e}{\bm{r}}^{e}_{t}\pm\bm{\epsilon^{\prime}}\pm\bm{\xi^{\prime}(t)}$ (20)
, where ${\bm{r}}^{i}_{t+1}=(f({\bm{x}}^{j}_{i},\Theta_{t})-y^{j}_{i})_{j=1}^{m},{\bm{r}}^{e}_{t}=(f({\bm{x}}^{j},\Theta_{t})-y^{j})$. ∎
@@
Further, from the derived expression equation 19, we observe that for different ${\bm{r}}^{i}_{t+1}$ we can replace ${\bm{r}}^{e}_{t}$ with ${\bm{r}}^{i}_{t}$ as the the pairwise cancellation of some error terms.
Appendix B Ablation Studies of Sampling Strategy
@@
As discussed in Sec. 3.3, we conduct ablation studies about the effect of sampling points and the size $p$ of sampling group. All experiments are conducted on the first three images of Kodak 24 dataset, and the average PSNR values are reported for reference.
B.1 The effect of Sampling points
@@
In this subsection, we examine sampling point selection. We randomly sample one point from each group in each iteration, referring to the resulting MLPs as randomly sampling (RI). Additionally, we implement a variant where random sampling occurs only at the beginning, denoting these MLPs as randomly sampling initially (RSI). We also introduce sampling based on the largest residual, termed SLR. Table 4 presents the average PSNR results while maintaining the same settings as in Experiment 6.1, demonstrating that our method significantly improves performance across various sampling strategies, with SLR yielding the best average performance. Consequently, we adopt SLR for sampling strategy.
@@
B.2 Sampling group and Training time
The size $p$ of the sampling group influences the level of data similarity within the group and hardware requirements. Intuitively, a smaller sampling group size reduces errors in estimating training dynamics and allows for more accurate tailored impacts on spectral bias, leading to improved representation performance. However, this leads to a larger matrix $\tilde{{\bm{K}}}$, which results in longer run times and greater memory usage.
To further explore the effect of different sizes of sampling groups, we follow the MLP architecture in Experiment 6.1. All training settings are remained. We vary the size from $16\times 16$ to $128\times 128$. Larger sizes correspond to coarser estimates of the dynamics. We report the average training time of each iteration.
As shown in Table 5, our method under all sampling sizes has the significant improvement comparing to the baseline models. Particularly, when the size is $128\times 128$ that means that the size of $\tilde{{\bm{K}}}_{e}$ is only $24\times 24$, it still improves the representation performance. That means that relatively low-precision estimates are sufficient to achieve representation improvements. Further, as the sampling group size decreases, representation accuracy improves; however, the extent of this improvement gradually diminishes with smaller group size. Specifically, reducing the group size from $64^{2}$ to $32^{2}$ nearly doubles the average increment $\Delta_{\text{avg}}$, while the increment becomes negligible when reducing from $32^{2}$ to $16^{2}$, while the time required for $16^{2}$ increases fivefold. This indicates that, in practical applications, there is no need to consider more precise sampling intervals; a relatively coarse partition suffices to achieve substantial improvements while maintaining efficiency.
@@ B.2 Sampling group and Training time
Appendix C Empirical Analysis on Simple Function Approximation
In this section, we first visualize more experimental results of Experiment 1 in our paper. Subsequently, we examine impacts on spectral bias of ReLU and SIREN with varying numbers of balanced eigenvalues under different group sizes from Experiment 2.
As can be seen in Fig. 6, the same trend as our observations in our paper, with the width increases, the differences on MSE curves and spectrum between gradient adjustments by ${\bm{S}}$, $\tilde{{\bm{S}}}$ and $\tilde{{\bm{S}}}_{e}$ gradually diminish when $end=13,15$. This is consistent with our Theorem 4.1 and the analysis of our Theorem 4.2. As shown in Fig. 7, 8 and 9, impacts on spectral bias are gradually amplified by increasing the number of balanced eigenvalues of $\tilde{{\bm{S}}}_{e}$, which is also consistent with the observations in our paper. The above empirical results corroborate our theoretical results and show that our inductive gradient adjustment (IGA) method amplify impacts on spectral bias by increasing the number of balanced eigenvalues.
Figure 6: Evolution of approximation error with training iterations on time domain and Fourier domain. The shaded area indicates training fluctuations, represented by twice the standard deviation. Line plots visualize the MSE loss curves of MLPs. Heatmaps show the relative error $\Delta_{k}$ on four frequency bands. SGD-$\tilde{{\bm{S}}}_{e}$-11 denotes that the MLP with ReLU optimized using gradients adjusted by $\tilde{{\bm{S}}}_{e}$ with $end=11$ and group size $p=8$.
Figure 7: Progressively amplified impacts on spectral bias of ReLU and SIREN by increasing the number of balanced eigenvalues of $\tilde{{\bm{S}}}_{e}$ when the group size is $8$. ReLU denotes that the MLP with ReLU optimized using vanilla gradients; ReLU-$\tilde{{\bm{S}}}_{e}$-2-IGA-8 denotes that the MLP with ReLU optimized using gradients adjusted by $\tilde{{\bm{S}}}_{e}$ with $end=2$ and group size is 8. Details of experimental setting can be found in Sec. 5.
Figure 8: Comparison of group sizes with varying balanced eigenvalues on Relative Error at 5, 10, 15, 20, 25, 30Hz of ReLU. IGA-1 denotes that the group size is 1, i.e., the eNTK-based gradient adjustment; IGA-4 denotes that the group size is 4 for IGA; Baseline denotes MLPs with ReLU activation function optimized by vanilla gradients. he shaded area indicates training fluctuations, represented by twice the standard deviation. Details of experimental setting can be found in Sec. 5.
Figure 9: Comparison of group sizes with varying balanced eigenvalues on Relative Error at 10, 20, 30, 40, 50, 60Hz of SIREN. IGA-1 denotes that the group size is 1, i.e., the eNTK-based gradient adjustment; IGA-4 denotes that the group size is 4 for IGA, i.e., inductive gradient adjustment; Baseline denotes the SIREN optimized by vanilla gradients. he shaded area indicates training fluctuations, represented by twice the standard deviation. Details of experimental setting can be found in Sec. 5.
@@ Appendix C Empirical Analysis on Simple Function Approximation
Appendix D per-scene results of 2D COLOR IMAGE APPROXIMATION
In this section, we provide metrics of each image and give more visualization. Peak signal-to-noise ratio (PSNR) results for 2D color image approximation are shown in Table 6. Structural Similarity Index Measure (SSIM) results for 2D color image approximation are shown in Table 7. Learned Perceptual Image Patch Similarity (LPIPS) by the “alex” in Zhang et al. (2018) of 2D color image approximation results are also reported in Table 8. We visualize the representation results of all methods on Kodak 5 and Kodak 6 in Fig. 10.
As can be seen in Table 6, 7 and 8, our IGA method achieves the best average results in three metrics across three model architectures. And other representation results can be found in Fig. 10.
@@
Figure 10: Visualization of 2D color image approximation results and PSNR values on Kodak 5 (top) and Kodak 6 (bottom). ReLU+IGA denotes that the MLPs with ReLU activation function are optimized by our IGA method. Details can be found in Sec. 6.1

### v2

@@ A.2 Basic Lemmas
where $\Delta f({\bm{x}}_{i},\Theta_{t})=({\bm{r}}_{t+1}-{\bm{r}}_{t})_{i}$ denotes the dynamics of $f({\bm{x}}_{i},\Theta)$ at time step $t$; $R_{0},\kappa$ are constants; $\nabla_{\Theta_{t}}f({\bm{x}}^{j_{i}},\Theta)^{\top}\nabla_{\Theta_{t}}f({\bm{x}}^{j},\Theta)$ is the entry in ${\tilde{{\bm{K}}}_{e}}(i,j)$.
@@
From the equation 17 of lemma A.4, we consider the following error:
@@
$\leq\epsilon\|\nabla_{\Theta_{t}}f({\bm{X}},\Theta_{t}){\bm{I}}{\bm{r}}_{t}\|\leq\epsilon\frac{\kappa R_{0}}{\sqrt{m}}.$
@@
We define that $\epsilon^{\prime}=A+B$. By the lemma A.4, we have that:
@@
${\bm{r}}^{i}_{t+1}=p\tilde{{\bm{K}}}_{e}{\bm{r}}^{e}_{t}\pm\bm{\epsilon^{\prime}}\pm\bm{\xi^{\prime}(t)},$ (19)
where ${\bm{r}}^{i}_{t+1}=(f({\bm{x}}^{j}_{i},\Theta_{t})-y^{j}_{i})_{j=1}^{m},{\bm{r}}^{e}_{t}=(f({\bm{x}}^{j},\Theta_{t})-y^{j})$. ∎
@@
Further, from the derived expression equation 18, we observe that for different ${\bm{r}}^{i}_{t+1}$ we can replace ${\bm{r}}^{e}_{t}$ with ${\bm{r}}^{i}_{t}$ as the pairwise cancellation of some error terms.
Appendix B Empirical analysis on simple function approximation
In this section, we first visualize more results of the experiment 1 in our paper. Subsequently, we illustrate impacts of IGA on spectral bias of ReLU and SIREN with varying numbers of balanced eigenvalues under varying group size $p$ from the experiment 2. As can be seen in Fig. 9, the same trend as our observations in our paper, with the width increases, the differences on MSE curves and spectrum between gradient adjustments by ${\bm{S}}$, $\tilde{{\bm{S}}}$ and $\tilde{{\bm{S}}}_{e}$ gradually diminish when $end=12,14$. This is consistent with our Theorem 3.1 and the analysis of our Theorem 3.2. As shown in Fig. 10, 11 and 12, impacts on spectral bias are gradually amplified by increasing the number of balanced eigenvalues of $\tilde{{\bm{S}}}_{e}$, which is also consistent with the observations in our paper. The above empirical results corroborate our theoretical results and show that our IGA method provides a chance to amplify impacts on spectral bias by increasing the number of balanced eigenvalues under varying group size $p$.
Appendix C Visual illustration for the sampling strategy of IGA
In this section, we aim to provide a visual illustration of the sampling strategy introduced in Sec. 3.4. As introduced in the main text, for 1D and 2D inputs, we sample data with the largest residuals in each non-overlapping interval and patch, respectively. For higher-dimensional settings like $N$ $d$-dimensional input coordinates ($d\geq 3$), represented as a tensor of shape $(n_{1},\dots,n_{d},d)$, we flatten the first $d$ dimensions into a 2D tensor ${\bm{X}}^{\prime}$ of shape $(N,d)$, where $N=n_{1}\times\dots\times n_{d}$ (which is evident), and then group it along the first dimension of ${\bm{X}}^{\prime}$. Then, the input with the largest residuals in each group are selected to form the sampled data ${\bm{X}}_{e}$. We take $d=3$ as an example to further illustrate our sampling strategy. Considering a 3D “Thai” statue, we first sample data over a $512\times 512\times 512$ volume grid with each voxel within the “Thai” volume assigned a $1$, and voxel outside the “Thai” volume assigned a 00 following Saragadam et al. (2023); Shi et al. (2024a). The 3D coordinates $\{x_{1},\dots,x_{n_{1}}\}\times\{y_{1},\dots,y_{n_{2}}\}\times\{z_{1},\dots,z_{n_{3}}\}$ of voxels defined by the “Thai” volume are the inputs of INRs. We denote these coordnates as ${\bm{X}}$. Generally, ${\bm{X}}$ is represented as tensor of shape $(n_{1},n_{2},n_{3},3)$ and $n_{1},n_{2},n_{3}$ depends on the resolution of the aforementioned volume grid and the object volume. As demonstrated in Fig. 6, we visualize ${\bm{X}}$ as a cube with dimensions $n_{1}\times n_{2}\times n_{3}$, composed of $N$ cubical elements. Each cubical element contains its corresponding coordinate. For example, if we define the $n_{1}$-dimension as the $x$-axis, the $n_{2}$-dimension as the $y$-axis, and the $n_{3}$-dimension as the $z$-axis (with positive directions indicated in Fig. 6), the red-marked cubical element contains the corresponding coordinate $(x_{n_{1}},y_{1},z_{n_{3}})$. Then, we use the flatten operation to unfold the cube sequentially along the dimensions, obtaining ${\bm{X}}^{\prime}$. Note that the flatten operation preserves adjacency. Therefore, we directly group ${\bm{X}}^{\prime}$ and sample one element from each group to form ${\bm{X}}_{e}$, marked green in Fig. 6.
This sampling strategy is simple yet effective. The adjacency during grouping helps obtain a smaller consistency bound $\epsilon$, thereby reducing the estimate error. Meanwhile, the flatten operation maintains adjacency while reducing computational overhead. Our extensive experimental results validate the effectiveness of this strategy. Note that there might be more sophisticated strategies via clustering or segmentation, which diverge from the primary focus of our work. We leave this for future research. The ablation study about hyperparameters $n$ and $p$ of the strategy are provided in Sec. F.
Figure 6: Visual illustration for the sampling strategy of IGA. The detailed description of the above diagram is in Sec. C.
Appendix D Discussion about IGA in Adam
In the current field of deep learning, Adam (Kingma, 2014) as a variant of vanilla gradient descent has been proven to be one of the most effective methods empirically (Deng et al., 2009; He et al., 2016; Vaswani, 2017) and is widely adopted by current INR models (Sitzmann et al., 2020; Mildenhall et al., 2021; Saragadam et al., 2023; Zhu et al., 2024; Shi et al., 2024a). However, due to the introduction of adaptive learning rates and momentum, the characterization of training dynamics like Eq. 17 induced by Adam becomes exceptionally challenging and remains an open problem. Intuitively, momentum is the linear combination of previous gradients, which implies that momentum has the similar direction with current adjusted gradients. This suggests that momentum is inherently compatible with our IGA. Therefore, we adjust gradients first by IGA and then update the momentum and learning rates by traditional Adam. Adaptive learning rates typically result in larger update steps for parameters (Kingma, 2014; Wilson et al., 2017; Reddi et al., 2019). For better convergence, we utilize $\lambda_{end+1}$ to balance eigenvalues in equation 5. Our extensive empirical studies in INRs (Table 8, 12, 14 in our paper) demonstrate the compatibility of IGA with the Adam optimizer.
Appendix E IGA on more activations
In the main text, we have evaluated our IGA on ReLU, PE (Tancik et al., 2020) and SIREN (Sitzmann et al., 2020). In this section, we apply our IGA to the most recent, meticulously crafted state-of-the-art activations with periodic variable, i.e., Gauss (Ramasinghe & Lucey, 2022), WIRE (Saragadam et al., 2023) and FINER (Zhu et al., 2024). For these activations, we follow their initial learning rates by their official guidelines. The same training strategy is adopted for these experiments. We set $end=20$ for all these activations, which are consistent with PE and SIREN. These experiments are conducted on the first 8 images of Kodak 24. We report the average PSNR in the Table. 5. Similar to our results in the main text, our IGA consistently improves the approximation accuracy for the evaluated activations.
Table 5: Peak signal-to-noise ratio (PSNR $\uparrow$) of 2D color image approximation results on more state-of-the-art activations with periodic variable. “Gauss” and “WIRE” denote that MLPs of Gaussian and Gabor Wavelet activate function by periodic variable, respectively. The detailed settings can be found in Sec. 5.1 and Sec. E.
Average Gauss WIRE FINER
Metric Vanilla +IGA Vanilla +IGA Vanilla +IGA
PSNR $\uparrow$ 33.79 34.45 30.37 31.49 35.12 35.55
Appendix F Ablation study
@@
As discussed in Sec. 3.4, we conduct ablation studies about the effect of sampling data and the group size $p$. All experiments are conducted on the first three images of Kodak 24 dataset, and the average PSNR values are reported for reference.
F.1 The effect of Sampling data
@@
In this subsection, we examine sampling data selection. We randomly sample one data from each group in each iteration, referring to the resulting MLPs as randomly sampling (RI). Additionally, we implement a variant where random sampling occurs only at the beginning, denoting these MLPs as randomly sampling initially (RSI). We also introduce sampling based on the largest residual, termed SLR. Table 6 presents the average PSNR results while maintaining the same settings as in Experiment 5.1, demonstrating that our method significantly improves performance across various sampling strategies, with SLR yielding the best average performance. Consequently, we adopt SLR for sampling strategy.
@@
F.2 Group size $p$ and Training time
The size $p$ of the sampling group influences the level of data similarity within the group and hardware requirements. Intuitively, a smaller $p$ reduces errors in estimating training dynamics and allows for more accurate tailored impacts on spectral bias, leading to improved representation performance. However, this leads to a larger matrix $\tilde{{\bm{K}}}$, which results in longer run times and greater memory usage.
To further explore the effect of various $p$, we follow the MLP architecture in Experiment 5.1. All training settings are remained. We vary $p$ from $16^{2}$ to $128^{2}$. Larger sizes correspond to coarser estimates of the dynamics. We report the average training time of each iteration.
As shown in Table 7, our method under all sampling sizes has the significant improvement comparing to the baseline models. Particularly, when the size $p$ is $128^{2}$ that means that the size of $\tilde{{\bm{K}}}_{e}$ is only $24^{2}$, it still improves the representation performance. That means that relatively low-precision estimates are sufficient to achieve representation improvements. Further, as $p$ decreases, representation accuracy improves; however, the extent of this improvement gradually diminishes with smaller group size $p$. Specifically, reducing the group size $p$ from $64^{2}$ to $32^{2}$ nearly doubles the average increment $\Delta_{\text{avg}}$, while the increment becomes negligible when reducing $p$ from $32^{2}$ to $16^{2}$, while the time required for $p=16^{2}$ increases fivefold. This indicates that, in practical applications, there is no need to consider more precise sampling intervals; a relatively coarse partition suffices to achieve substantial improvements while maintaining efficiency.
@@ F.2 Group size $p$ and Training time
Appendix G Experimental details and per-scene results of 2D color image approximation
In this section, we first include additional experimental details that were omitted in the main text due to space constraints. Then we provide the detailed per-scene results for all four metrics and more visualization results.
As illustrated in the main text, we follow the learning rate schedule of Shi et al. (2024a), which maintains a fixed rate for the first $3$K iterations and then reduces it by $0.1$ for another $7$K iterations. For IGA, we set initial learning rates as $35e-$ 3 for ReLU activation and $31e-$ 3 for Sine activation. For all baseline models, we set initial learning rates as $31e-$ 3 due to poor performance observed with $35e-$ 3. Full-batch training is adopted.
We report four metrics to comprehensively evaluate the representation performance, i.e., peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), multi-scale structural similarity index measure (MS-SSIM) and learned perceptual image patch similarity (LPIPS), in Table 8-11. Specifically, PSNR is used to measure average representation accuracy based on MSE, while SSIM, MS-SSIM and LPIPS are more sensitive to the representation of high-frequency information. As shown in Table 8-11, all methods improve representation quality across the four metrics. However, the improvements from FR and BN are inconsistent. Although they enhance frequency-sensitive metrics to some extent, they degrade average representation accuracy and even lead to negative effects in certain cases. In contrast, our IGA method achieves consistent improvements across almost all scenarios and metrics, demonstrating the superiority of tailor impacts on spectral bias induced by our IGA. We further visualize more representation results in Fig. 7, which is consistent with our observation of Fig. 4 that mprovements of IGA are uniformly distributed across most frequency bands.
Figure 7: Visual examples of 2D color image approximation results by different training dynamics methods on Kodak 4. Enlarged views of the regions labeled by red boxes are provided. The residuals of these regions in the Fourier domain are visualized through heatmaps in the top line. The increase in error corresponds to the transition of colors in the heatmaps from blue to red. ReLU+IGA denotes that the MLPs with ReLU activations are optimized by our IGA method. Details can be found in Sec. 5.1 and Sec. G.
Appendix H Experimental details and per-scene results of 3D shape representation
In this section, we firstly introduce the detailed training strategy of the 3D shape representation experiment. Then we provide the intersection over union (IOU) and Chamfer distance metric values of each 3D shape and give more visualization.
Following previous works (Saragadam et al., 2023; Shi et al., 2024a; Cai et al., 2024), we train all models for 200 epochs with a learning rate decay exponentially to $0.1$ of initial rates. We set the initial learning rate as $32e-$ 3 for ReLU and PE; For SIREN, we set the initial learning rate as $45e-$ 4. FR, BN and our IGA method are adopted the same learning rate with baseline models. We typically set $end=7$ for ReLU and PE, and $end=14$ for SIREN. The per-scene results are reported in Table 12 and 13. It can be observed that our IGA method achieves improvements in IOU metrics across all objects, and achieves the best performance in average metrics of all objects. We visualize the Lucy object in Fig. 8.
@@ Appendix H Experimental details and per-scene results of 3D shape representation
Figure 8: Visualization of 3D shape representation results and IOU values on the Lucy. ReLU+IGA denotes that the MLPs with ReLU activations are optimized by our IGA method. Details can be found in Sec. 5.2 and Sec. H.
Appendix I Experimental details and per-scene results of Learning Neural Radiance Fields
In this section, we firstly introduce some training details. Then we provide metrics of each scenario and give more visualization. As previously discussed in Experiment 5.3, we apply our method on the original NeRF model (Mildenhall et al., 2021).
Specifically, given a ray $i$ from the camera into the scene, the MLP takes the 5D coordinates (3D spatial locations and 2D view directions) of $N$ points along the ray as inputs and outputs the corresponding color $\mathbf{c}$ and volume density $\sigma$. Then $\mathbf{c}$ and $\sigma$ are combined using numerical volume rendering to obtain the final pixel color of the ray $i$. We set each ray as one group. For each group, we sample the point with the maximum integral weight. The original NeRF adopts two models: “coarse” and “fine” models. We adjust the gradients of the “fine” model as it captures more high-frequency details and remain gradients of the “coarse” model.
The “NeRF-PyTorch” codebase (Yen-Chen, 2020) is used and we follow its default settings. Besides, we also compare to previous training dynamics methods, i.e., FR and BN. Their hyperparameters follow their publicly available codes. Generally, we set the $end=25$. For more complex scenes, such as the branches and leaves of the ficus, we set the $end=30$ to achieve better learning of high frequencies. We report per-scene results in Table 14 and visualize several scenes in Fig. 13.
Figure 9: Evolution of approximation error with training iterations on time domain and Fourier domain. The shaded area indicates training fluctuations, represented by twice the standard deviation. Line plots visualize the MSE loss curves of MLPs. Heatmaps show the relative error $\Delta_{k}$ on four frequency bands. SGD-$\tilde{{\bm{S}}}_{e}$-10 denotes that the MLP with ReLU optimized using gradients adjusted by $\tilde{{\bm{S}}}_{e}$ with $end=10$ and group size $p=8$. Please zoom in for better review.
Figure 10: Progressively amplified impacts on spectral bias of ReLU and SIREN by increasing the number of balanced eigenvalues of $\tilde{{\bm{S}}}_{e}$ when the group size $p$ is $8$. ReLU denotes that the MLP with ReLU optimized using vanilla gradients; ReLU-IGA-8-Balance-2 denotes that the MLP with ReLU optimized using gradients adjusted by $\tilde{{\bm{S}}}_{e}$ with $end=1$ (for Adam, there are two balanced eigenvalues) and group size $p$ is 8. Details of experimental setting can be found in Sec. 4. Please zoom in for better review.
Figure 11: Comparison of group sizes with varying balanced eigenvalues on Relative Error $\Delta_{k}$ at 5, 10, 15, 20, 25, 30Hz of ReLU. IGA-1 denotes that the group size $p$ is 1, i.e., $\tilde{{\bm{K}}}$-based gradient adjustment; IGA-4 denotes that the group size $p$ is 4 for IGA; Baseline denotes MLPs with ReLU activations optimized by vanilla gradients. The shaded area indicates training fluctuations, represented by twice the standard deviation. Details of experimental setting can be found in Sec. 4. Please zoom in for better review.
Figure 12: Comparison of group sizes with varying balanced eigenvalues on Relative Error $\Delta_{k}$ at 10, 20, 30, 40, 50, 60Hz of SIREN. IGA-1 denotes that the group size $p$ is 1, i.e., $\tilde{{\bm{K}}}$-based gradient adjustment; IGA-4 denotes that the group size $p$ is 4 for IGA, i.e., inductive gradient adjustment; Baseline denotes the SIREN optimized by vanilla gradients. The shaded area indicates training fluctuations, represented by twice the standard deviation. Details of experimental setting can be found in Sec. 4. Please zoom in for better review.